# Temporal Sparse Autoencoders: Leveraging the Sequential Nature of Language for Interpretability

### v1

@@ Abstract
Translating the internal representations and computations of models into concepts that humans can understand is a key goal of interpretability. While recent dictionary learning methods such as Sparse Autoencoders (SAEs) provide a promising route to discover human-interpretable features, they suffer from a variety of problems, including a systematic failure to capture the rich conceptual information that drives linguistic understanding. Instead, they exhibit a bias towards shallow, token-specific, or noisy features, such as “the phrase ‘The’ at the start of sentences”. In this work, we propose that this is due to a fundamental issue with how dictionary learning methods for LLMs are trained. Language itself has a rich, well-studied structure spanning syntax, semantics, and pragmatics; however, current unsupervised methods largely ignore this linguistic knowledge, leading to poor feature discovery that favors superficial patterns over meaningful concepts. We focus on a simple but important aspect of language: semantic content has long-range dependencies and tends to be smooth over a sequence, whereas syntactic information is much more local. Building on this insight, we introduce Temporal Sparse Autoencoders (T-SAEs), which incorporate a novel contrastive loss encouraging consistent activations of high-level features over adjacent tokens. This simple yet powerful modification enables SAEs to disentangle semantic from syntactic features in a self-supervised manner. Across multiple datasets and models, T-SAEs recover smoother, more coherent semantic concepts without sacrificing reconstruction quality. Strikingly, they exhibit clear semantic structure despite being trained without explicit semantic signal, offering a new pathway for unsupervised interpretability in language models.
footnotetext: $^{*},^{\dagger}$ Equal contribution.
@@ 1 Introduction
Interpretability aims to translate the internal representations and computations of language models into concepts that humans can understand, evaluate, and ultimately control. In practice, the most useful insights often involve high-level drivers of model behavior, such as the semantics a model encodes or the state it is operating in, rather than surface-level statistical patterns. Recent dictionary learning methods, such as Sparse Autoencoders (SAEs) [1], have shown promise in explaining language models. By projecting dense latent representations into a sparse, human-interpretable feature space, SAEs enable both the recovery of known linguistic patterns and the discovery of novel concepts within models.
@@
However, when applied to large language models (LLMs), SAEs frequently fall short of this goal. The features they recover are often token-specific, local, and noisy, capturing superficial syntactic patterns (e.g., “the phrase ‘The’ at the start of sentences" [^1] Neuronpedia, Feature 11795 of Gemmascope’s Gemma2-2b, Residual, 16k SAE or “Sentence endings or periods," Figure 1) rather than coherent, high-level semantic concepts. One interpretation of this phenomenon is that LLMs themselves fail to encode deep semantic structure, functioning instead as sophisticated next-token predictors. A more plausible explanation, however, is that current concept discovery methods are inadequate; their design biases them toward recovering shallow patterns even when richer structure exists in the underlying representations. We argue that this limitation stems from a fundamental issue with how SAEs are formulated. Human language is inherently structured: meaning is conveyed through context and semantics that evolve smoothly over time, while syntax is governed by more local dependencies [2]. Yet current dictionary learning methods ignore this sequential structure, treating tokens as independent and stripped of context.
@@
To address this gap, we introduce the notion of temporal consistency, or the property that high-level semantic features of a sequence remain stable over adjacent (or nearby) tokens, while low-level syntactic features may fluctuate more rapidly. For example, in the sentence “Photosynthesis is the process by which plants convert sunlight into energy,” a temporally consistent semantic feature might represent “discussion of plant biology" or “scientific explanation." This feature should remain active throughout the entire sentence, because the semantic content does not change from word to word. In contrast, syntactic features such as “capitalized first word” or “plural noun” activate only at specific tokens (e.g., “Photosynthesis,” “plants”), reflecting local rather than global structure.
@@
Building on this principle, we propose Temporal Sparse Autoencoders (T-SAEs), a simple modification to SAEs that incorporates a temporal contrastive loss encouraging high-level features to activate consistently over adjacent tokens (Figure 1). This modification, grounded in linguistic intuition, enables T-SAEs to more reliably capture semantic features while still disentangling them from low-level syntactic ones. Despite being trained only on a self-supervised context-similarity objective, T-SAEs yield representations with significantly improved semantic structure, without requiring any explicit semantic signal. Through experiments, we show that T-SAEs consistently recover higher-level semantic and contextual concepts, exhibit smoother and more temporally consistent activations over sequences, and remain competitive with existing SAEs on standard benchmarks such as reconstruction quality.
@@
Figure 1: A) Human language production involves high-level features such as semantic content and surrounding context, as well as low-level features such as syntactical requirements and specific word choices. B) While existing SAEs mostly recover syntactic information, Temporal SAEs balance recovery of semantics, syntax, and context. C) When decomposing a sequence composed of three passages: Newton’s Principia, an MMLU genetics question, and the Bhagavat Gita, Temporal SAEs (bottom) are able to smoothly detect the semantic shifts in the passage, with highly active features strongly correlating to the true content of the text, whereas existing SAEs (such as Matryoshka, top), are much noisier, varying on almost a per-token basis, and do not easily depict these shifts.
Our contributions are the following:
@@
    We introduce a simple data-generating process for language that distinguishes between high-level, temporally consistent semantic variables and local, low-level syntactic variables. This framework formalizes how we expect language models to encode linguistic information and provides guidance for designing better interpretability methods.
@@
    We propose Temporal SAEs, which partition latent features into semantic and syntactic components. We introduce a novel temporal contrastive loss which enforces consistency of high-level activations across tokens, encouraging T-SAEs to disentangle semantic and syntactic features in a self-supervised manner.
@@
    Through experiments on multiple models and datasets, we show that T-SAEs: a) Recover semantic and contextual information more reliably than state-of-the-art SAEs, b) Exhibit improved disentanglement between high- and low-level features, c) Maintain competitive performance on standard reconstruction benchmarks, and d) Provide practical interpretability benefits, including a case study on human preference data understanding and model steering.
@@ 2 Related Work
In recent years, SAEs [3] have emerged as a popular mechanistic interpretability technique for self-supervised concept discovery [4, 5, 6, 7, 8]. While they were initially promising for addressing the problem of polysemanticity, where a single neuron in a model can represent multiple concepts [9], in practice, they have been shown to create new problems, such as feature splitting and absorption [10], where concepts are split across multiple features or absorbed into less interpretable sub-features. To address these subsequent issues, methods such as Matryoshka SAEs [11] and Transcoders [12] have been proposed, which learn hierarchical and causal features. Recent work has also proposed learning dictionary features that are constrained to the data manifold [13] and reflect intuition about the geometry of model latent spaces [14], allowing for the recovery of heterogeneous concepts. However, all of these works assume a fully unsupervised objective for learning SAEs, treating each token in the training data as i.i.d., without acknowledging the temporal aspect of language and other sequential modalities. As a result, SAEs are known to suffer from a variety of problems, including “dense" activation behavior [15] and lack of utility for steering [16, 17].
@@
Linguistics, Cognitive Science, and Neuroscience. Many foundational works in linguistics study the relationship between syntactic structure and semantic meaning [2], with nearly all major theories recognizing a fundamental distinction between semantics and syntax. Evidence of the two having distinct representations has been found in developmental psychology [18] and neuroscience [19, 20]. In computational linguistics, separate approaches emerged to model language purely syntactically via Hidden Markov Models [21] or through a bag-of-words approach to model topics with Latent Dirichlet Allocation [22]. Griffiths et al. [23] combine the two into a single model consisting of an HMM where one “semantic" class denotes a topic model which samples words in an LDA-like fashion. Importantly, Griffiths et al. [23] argue that semantics in language exhibit long-range behavior, with different words or sentences in the same document having similar semantic content, whereas syntax is mostly dependent on short-range interactions, motivating our method.
@@
Dictionary learning with natural priors. The seminal work of Olshausen and Field [24] found that decomposing natural images in a sparse linear manner leads to Gabor-like receptive filters, similar to what is found in the visual cortex without any priors on visual data, ushering in work on sparse coding through dictionary learning [25, 26]. Later approaches then realized the benefit of taking data priors into the dictionary learning process. For instance, low-rankness has proved an effective parsimonious prior beyond sparsity [27, 28], multi-scale structure in medical imaging [29], and Luo et al. [30] argued that sequential frames in videos exhibit temporal consistency. Our work draws parallels to this trajectory of research in compressive sensing and dictionary learning by introducing structural priors to the unsupervised learning of SAEs.
@@ 3 Our Framework: Temporal Sparse Autoencoders
In this section, we describe our data generating process and the assumptions behind our proposed method. Then, we propose a novel training approach for Temporal SAEs to achieve temporal consistency in practice.
@@ 3.1 Formulating the Data Generating Process
The mapping $g$ can be decomposed as:
$g(\mathbf{h}_{t},\mathbf{l}_{t})=g_{H}(\mathbf{h}_{t})+g_{L}(\mathbf{l}_{t})$ (1)
where $g_{H}:\mathcal{H}\rightarrow\mathbb{R}^{d}$ captures high-level structure and while $g_{L}:\mathcal{L}\rightarrow\mathbb{R}^{d}$ encodes residual low-level information. Additionally, we have that
$\|g(\mathbf{h}_{t},\mathbf{l}_{t})-\mathbf{x}_{t}\|_{2}\leq\|g_{H}(\mathbf{h}_{t})-\mathbf{x}_{t}\|_{2}\leq\epsilon.$ (2)
In other words, $\mathbf{h}_{t}$ can reconstruct $\mathbf{x}_{t}$, but $\mathbf{l}_{t}$ encodes residual low-level information orthogonal to the $\mathbf{h}_{t}$ that aids in reconstruction.
@@ 3.2 Temporal Sparse Autoencoders
$\mathbf{f}(\mathbf{x}_{t})$ $=\sigma(\mathbf{W}^{\mathrm{enc}}\mathbf{x}_{t}+\mathbf{b}^{\mathrm{enc}}),$
$\hat{x}(\mathbf{f})$ $=\mathbf{W}^{\mathrm{dec}}\mathbf{f}(\mathbf{x}_{t})+\mathbf{b}^{\mathrm{dec}}.$
Here, $\mathbf{W}^{\mathrm{enc}}\in\mathbb{R}^{m\times d}$ is the encoder matrix, and $\mathbf{W}^{\mathrm{dec}}\in\mathbb{R}^{d\times m}$ is the decoder comprised of high-level features $\mathbf{W}^{\mathrm{dec}}_{0:h}\in\mathbb{R}^{d\times h}$ and low-level features $\mathbf{W}^{\mathrm{dec}}_{h:m}\in\mathbb{R}^{d\times(m-h)}$ such that their concatenation equals $\mathbf{W}^{\mathrm{dec}}$. The encoder and decoder bias are $\mathbf{b}^{\mathrm{enc}}\in\mathbb{R}^{d}$ and $\mathbf{b}^{\mathrm{dec}}\in\mathbb{R}^{d}$, respectively. We define the following loss function, where the high-level features $\mathbf{f}_{0:h}(\mathbf{x}_{t})$ should reconstruct the input and the low-level features $\mathbf{f}_{h:m}(\mathbf{x}_{t})$ should reconstruct the residual, as discussed in Assumption 2, similar to the Matryoshka SAE objective in [11].
@@
$\mathcal{L}_{H}$ $=\|\mathbf{x}_{t}-\mathbf{W}^{\mathrm{dec}}_{0:h}\mathbf{f}_{0:h}(\mathbf{x}_{t})+\mathbf{b}^{\mathrm{dec}}\|^{2}_{2},$
$\mathcal{L}_{L}$ $=\|\mathbf{x}_{t}-\mathbf{W}^{\mathrm{dec}}\mathbf{f}(\mathbf{x}_{t})+\mathbf{b}^{\mathrm{dec}}\|^{2}_{2}.$
We then add a training objective that encourages $\mathbf{W}^{\mathrm{enc}}_{0:h}$ to learn temporally-consistent features following Assumption 1 about $\mathbf{h}_{t}$: high-level features should be similar for two tokens from the same sequence, particularly for two adjacent tokens. To enforce this, we add a contrastive term to the loss function that encourages $\mathbf{W}^{\mathrm{enc}}_{0:h}\mathbf{x}_{t}$ to be similar to $\mathbf{W}^{\mathrm{enc}}_{0:h}\mathbf{x}_{t-1}$. Let $\mathbf{z}_{t}$ be the high-level features $\mathbf{f}_{0:h}(\mathbf{x}_{t})$, and let $s(\mathbf{x},\mathbf{y})$ be the cosine similarity between vectors $\mathbf{x}$ and $\mathbf{y}$ in the same latent space. Our full loss over a batch is subsequently
@@
$\mathcal{L}_{\mathrm{contr}}$ $=-\frac{1}{N}\sum_{i=1}^{N}\log\frac{e^{s(\textbf{z}_{t}^{(i)},\textbf{z}_{t-1}^{(i)})}}{\sum_{j=1}^{N}e^{s(\textbf{z}_{t}^{(i)},\textbf{z}_{t-1}^{(j)})}}-\frac{1}{N}\sum_{j=1}^{N}\log\frac{e^{s(\textbf{z}_{t-1}^{(j)},\textbf{z}_{t}^{(j)})}}{\sum_{i=1}^{N}e^{s(\textbf{z}_{t-1}^{(i)},\textbf{z}_{t}^{(j)})}},$
@@
While the contrastive loss is applied only to high-level features, for low-level, token-specific features, we do not apply any constraints. By nature of fitting the residual data left unexplained by the high-level component of the network, our loss naturally encourages the low-level latents to capture remaining, fluctuating features over a sequence. We note that this contrastive loss can be applied to any SAE training scheme; we focus on the Matryoshka objective because it naturally partitions the feature space.
Figure 2: TSNE visualizations of Pythia-160m SAE decompositions of MMLU questions, labeled by question category (left), question number (middle column), and token part-of-speech (right). We see that the high-level features from T-SAEs (top) recover semantic and contextual information. The low-level features of T-SAEs (middle row), as well as Matryoshka SAEs (bottom), recover syntactic information.

### v2

@@ Abstract
Translating the internal representations and computations of models into concepts that humans can understand is a key goal of interpretability. While recent dictionary learning methods such as Sparse Autoencoders (SAEs) provide a promising route to discover human-interpretable features, they often only recover token-specific, noisy, or highly local concepts. We argue that this limitation stems from neglecting the temporal structure of language, where semantic content typically evolves smoothly over sequences. Building on this insight, we introduce Temporal Sparse Autoencoders (T-SAEs), which incorporate a novel contrastive loss encouraging consistent activations of high-level features over adjacent tokens. This simple yet powerful modification enables SAEs to disentangle semantic from syntactic features in a self-supervised manner. Across multiple datasets and models, T-SAEs recover smoother, more coherent semantic concepts without sacrificing reconstruction quality. Strikingly, they exhibit clear semantic structure despite being trained without explicit semantic signal, offering a new pathway for unsupervised interpretability in language models.
@@ 1 Introduction
Interpretability aims to translate the internal representations and computations of language models into concepts that humans can understand Kim et al. (2018), evaluate (Koh et al., 2020), and ultimately control (Bhalla et al., 2025). In practice, the most useful insights involve high-level drivers of model behavior, such as the semantics a model encodes or the state it is operating in, rather than surface-level statistical patterns. Recent dictionary learning methods, such as Sparse Autoencoders (SAEs), have shown promise in explaining language models (Bricken et al., 2023). By projecting dense latent representations into a sparse, human-interpretable feature space, SAEs enable both the recovery of known linguistic patterns and the discovery of novel concepts within models.
@@
However, when applied to large language models (LLMs), SAEs frequently fall short of this goal. The features they recover are often token-specific, local, unstable, and noisy, capturing superficial syntactic patterns (e.g., “the phrase ‘The’ at the start of sentences [^1] Neuronpedia, Feature 11795 of Gemmascope’s Gemma2-2b 16k SAE, see Appendix D for more examples. or “Sentence endings or periods,” Figure 1) rather than coherent, high-level semantic concepts (Menon et al., 2025; Paulo and Belrose, 2025). One interpretation of this phenomenon is that LLMs themselves fail to encode deep semantic structure, functioning instead as sophisticated next-token predictors. A more plausible explanation, however, is that current concept discovery methods are inadequate; their design biases them toward recovering shallow patterns even when richer structure exists in the underlying representations. We argue that this limitation stems from a fundamental issue with how SAEs are formulated. Human language is inherently structured: meaning is conveyed through context and semantics that evolve smoothly over time, while syntax is governed by more local dependencies (Chomsky, 1965; Griffiths et al., 2004). Yet current dictionary learning methods ignore this sequential structure, treating tokens as independent and stripped of context.
@@
To address this gap, we introduce the notion of temporal consistency, or the property that high-level semantic features of a sequence remain stable over adjacent (or nearby) tokens, while low-level syntactic features may fluctuate more rapidly. For example, in the sentence “Photosynthesis is the process by which plants convert sunlight into energy,” a temporally consistent semantic feature might represent “discussion of plant biology” or “scientific explanation.” We expect this feature to be active over the course of the sequence rather than simply on a few specific tokens. In contrast, syntactic features such as “capitalized first word” or “plural noun” activate only at specific tokens (e.g., “Photosynthesis,” “plants”), reflecting local rather than global structure.
@@
Building on this principle, we propose Temporal Sparse Autoencoders (T-SAEs), a simple modification to SAEs that incorporates a temporal contrastive loss encouraging high-level features to activate consistently over adjacent tokens (Figure 1). This modification, grounded in linguistic intuition, enables T-SAEs to more reliably capture semantic features while still disentangling them from low-level syntactic ones. Despite being trained only on a self-supervised context-similarity objective, T-SAEs yield representations with significantly improved semantic structure, without requiring any explicit semantic signal. Through experiments, we show that T-SAEs consistently recover higher-level semantic and contextual concepts, exhibit smoother and more temporally consistent activations over sequences, and remain competitive with existing SAEs on standard benchmarks such as reconstruction quality. Our contributions are the following:
@@
Figure 1: A) Human language production involves high-level features such as semantic content and surrounding context, as well as low-level features such as syntactical requirements and specific word choices. B) While existing SAEs mostly recover syntactic information, T-SAEs balance recovery of semantics, syntax, and context. C) When decomposing a sequence composed of three passages: Newton’s Principia, an MMLU genetics question, and the Bhagavat Gita (in English), T-SAEs (bottom) smoothly detect the semantic shifts in the passage, with highly active features strongly correlating to the true content of the text, whereas existing SAEs (such as Matryoshka, top), are much noisier, varying on almost a per-token basis, and do not easily depict these shifts.
@@
    We introduce a simple data-generating process for language that distinguishes between high-level, temporally consistent semantic variables and low-level, local syntactic variables. This framework formalizes how we expect language models to encode linguistic information and provides guidance for designing better interpretability methods.
@@
    Building on this framework, we propose Temporal SAEs, which partition latent features into semantic and syntactic components. We introduce a novel temporal contrastive loss which enforces consistency of high-level activations across sequences, encouraging T-SAEs to disentangle semantic and syntactic features in a self-supervised manner.
@@
    Through experiments on multiple models and datasets, we show that T-SAEs: a) Recover semantic and contextual information more reliably than existing SAEs, b) Exhibit improved disentanglement between high- and low-level features, c) Maintain competitive performance on standard reconstruction benchmarks, and d) Provide practical interpretability benefits, including a case study on safety-related concepts.
@@ 2 Related Work
In recent years, SAEs have emerged as a popular mechanistic interpretability technique for unsupervised concept discovery (Bricken et al., 2023; Templeton et al., 2024; Gao et al., 2024).While they were initially promising for addressing the problem of polysemanticity, where a single neuron in a model can represent multiple concepts (Elhage et al., 2022), in practice, they have been shown to create new problems, such as feature splitting and absorption (Chanin et al., 2024), where concepts are split across multiple features or absorbed into less interpretable sub-features. To address these subsequent issues, methods such as Matryoshka SAEs (Bussmann et al., 2025) and Transcoders (Paulo et al., 2025) have been proposed, which learn hierarchical and causal features. Recent work has also proposed learning dictionary features that are constrained to the data manifold (Fel et al., 2025) and reflect intuition about the geometry of model latent spaces (Hindupur et al., 2025), allowing for the recovery of heterogeneous concepts. However, all of these works assume a fully unsupervised objective for learning SAEs, treating each token in the training data as i.i.d., without acknowledging the temporal aspect of language and other sequential modalities. As a result, SAEs are known to suffer from a variety of problems, including “dense” activation behavior (Sun et al., 2025) and lack of utility for steering (Bhalla et al., 2025; Wu et al., 2025). The works most relevant to ours use temporal signal as weak supervision to learn more causal SAE features Menon et al. (2025) and address absorption Li and Ren (2025).
@@
Linguistics, Cognitive Science, and Neuroscience. Many foundational works in linguistics study the relationship between syntactic structure and semantic meaning (Chomsky, 1965), with nearly all major theories recognizing a fundamental distinction between semantics and syntax. Evidence of the two having distinct representations has been found in developmental psychology (Brown, 1973) and neuroscience (Neville et al., 1992; Zhang et al., 2025). In computational linguistics, separate approaches emerged to model language purely syntactically via Hidden Markov Models (Manning and Schutze, 1999) or through a bag-of-words approach to model topics with Latent Dirichlet Allocation (Blei et al., 2003). Griffiths et al. (2004) combine the two into a single model consisting of an HMM where one “semantic” class denotes a topic model which samples words in an LDA-like fashion. Importantly, Griffiths et al. (2004) argue that semantics in language exhibit long-range behavior, with different words or sentences in the same document having similar semantic content, whereas syntax is mostly dependent on short-range interactions, motivating our method.
@@
Dictionary and representation learning with natural priors. Olshausen and Field (1996) found that decomposing natural images in a sparse linear manner leads to Gabor-like receptive filters, similar to what is found in the visual cortex without any priors on visual data, ushering in work on sparse coding through dictionary learning (Tošić and Frossard, 2011; Dumitrescu and Irofti, 2018). Later approaches then realized the benefit of taking data priors into the dictionary learning process, such as low-rank (Davenport and Romberg, 2016; Vu and Monga, 2017) or multi-scale structure for images (Ong and Lustig, 2016), and temporal smoothness for video (Luo et al., 2019). Early works in feature and representation learning used time-contrastive approaches, assuming smoothness (Oord et al., 2018) for predictive coding, and nonstationarity (Hyvarinen and Morioka, 2016) and sparse transitions (Klindt et al., 2020) for disentanglement. Our work draws parallels to this trajectory of research in dictionary learning and representation learning by introducing structural priors to the unsupervised learning of SAEs.
@@ 3.1 Formulating the Data Generating Process
The mapping $g$ is hierarchical in the sense that it satisfies $0=\|g(\mathbf{h}_{t},\mathbf{l}_{t})-\mathbf{x}_{t}\|\leq\|g(\mathbf{h}_{t},\mathbf{0})-\mathbf{x}_{t}\|\leq\epsilon$. In other words, $\mathbf{h}_{t}$ can reconstruct $\mathbf{x}_{t}$ up to $\epsilon$, but $\mathbf{l}_{t}$ contains signal about $\mathbf{x}_{t}$ that is not explained by $\mathbf{h}_{t}$.
@@ 3.2 Temporal Sparse Autoencoders
$\mathbf{f}(\mathbf{x}_{t})=\sigma(\mathbf{W}^{\mathrm{enc}}\mathbf{x}_{t}+\mathbf{b}^{\mathrm{enc}}),$ $\hat{x}(\mathbf{f})=\mathbf{W}^{\mathrm{dec}}\mathbf{f}(\mathbf{x}_{t})+\mathbf{b}^{\mathrm{dec}}.$
Here, $\mathbf{W}^{\mathrm{enc}}\in\mathbb{R}^{m\times d}$ is the encoder matrix, and $\mathbf{W}^{\mathrm{dec}}\in\mathbb{R}^{d\times m}$ is the decoder comprised of high-level features $\mathbf{W}^{\mathrm{dec}}_{0:h}\in\mathbb{R}^{d\times h}$ and low-level features $\mathbf{W}^{\mathrm{dec}}_{h:m}\in\mathbb{R}^{d\times(m-h)}$ such that their concatenation equals $\mathbf{W}^{\mathrm{dec}}$. The encoder and decoder bias are $\mathbf{b}^{\mathrm{enc}}\in\mathbb{R}^{d}$ and $\mathbf{b}^{\mathrm{dec}}\in\mathbb{R}^{d}$, respectively. We define the following loss function, where the high-level features $\mathbf{f}_{0:h}(\mathbf{x}_{t})$ should reconstruct the input and the low-level features $\mathbf{f}_{h:m}(\mathbf{x}_{t})$ should reconstruct the residual, as discussed in Assumption 2, similar to the Matryoshka SAE objective in (Bussmann et al., 2025).
@@
$\mathcal{L}_{H}=\|\mathbf{x}_{t}-\mathbf{W}^{\mathrm{dec}}_{0:h}\mathbf{f}_{0:h}(\mathbf{x}_{t})+\mathbf{b}^{\mathrm{dec}}\|^{2}_{2},$ $\mathcal{L}_{L}=\|\mathbf{x}_{t}-\mathbf{W}^{\mathrm{dec}}\mathbf{f}(\mathbf{x}_{t})+\mathbf{b}^{\mathrm{dec}}\|^{2}_{2}.$
We then add a training objective that encourages $\mathbf{W}^{\mathrm{enc}}_{0:h}$ to learn temporally-consistent features following Assumption 1 about $\mathbf{h}_{t}$: high-level features should be similar for two tokens from the same sequence, particularly for two adjacent tokens. To enforce this, we add a contrastive term to the loss function that encourages $\mathbf{W}^{\mathrm{enc}}_{0:h}\mathbf{x}_{t}$ to be similar to $\mathbf{W}^{\mathrm{enc}}_{0:h}\mathbf{x}_{t-1}$. In addition to encouraging temporal consistency between pairs, discouraging similarity across different samples prevents smoothness collapse (where high-level features are constant) and encourages full use of the high-level feature space. Let $\mathbf{z}_{t}$ be the high-level features $\mathbf{f}_{0:h}(\mathbf{x}_{t})$, and let $s(\mathbf{x},\mathbf{y})$ be the cosine similarity between vectors $\mathbf{x}$ and $\mathbf{y}$ in the same latent space. Our full loss over a batch is subsequently
@@
$\mathcal{L}_{\mathrm{contr}}$ $=-\frac{1}{N}\sum_{i=1}^{N}\log\frac{\exp({s(\textbf{z}_{t}^{(i)},\textbf{z}_{t-1}^{(i)})})}{\sum_{j=1}^{N}\exp({s(\textbf{z}_{t}^{(i)},\textbf{z}_{t-1}^{(j)})})}-\frac{1}{N}\sum_{j=1}^{N}\log\frac{\exp({s(\textbf{z}_{t}^{(j)},\textbf{z}_{t-1}^{(j)})})}{\sum_{i=1}^{N}\exp({s(\textbf{z}_{t}^{(i)},\textbf{z}_{t-1}^{(j)})})},$
@@
While the contrastive loss is applied only to high-level features, for low-level, token-specific features, we do not apply any constraints. By nature of fitting the residual data left unexplained by the high-level component of the network, our loss naturally encourages the low-level latents to capture remaining, fluctuating features over a sequence.
Figure 2: TSNE visualizations of Pythia-160m SAE decompositions of MMLU questions, labeled by question category (left), question number (middle column), and token part-of-speech (right). We see that T-SAE high-level features (top) recover semantics and context. The low-level features of T-SAEs (middle row), as well as Matryoshka SAEs (bottom), recover syntactic information.